# Topological restrictions on relatively Anosov representations

### v1

@@ 2.5. Relatively Anosov subgroups
Let $(\Gamma,\mathcal{P})$ be a non-elementary relatively hyperbolic group which is virtually torsion-free. The limit set of a subgroup Γsuperscriptnormal-$\Gamma^{\prime}$ of $\Gamma$ is a singleton if and only if Γsuperscriptnormal-$\Gamma^{\prime}$ is virtually a peripheral subgroup.
@@
By passing to exterior powers we may assume that $k=1$, see [Zhu21, Def. 10.2] and [ZZ23, Ex. 13.2]). Since $\Gamma_{0}$ is non-elementary there exists $\gamma\in\Gamma$ such that $\rho(\gamma)$ is $1$-biproximal (see [Zhu21, Prop. 4.6] and [ZZ23, Prop. 4.2]), and $\xi_{\rho}^{1}(\gamma^{+})$ and $\xi_{\rho}^{d-1}(\gamma^{+})$ (resp. $\xi_{\rho}^{1}(\gamma^{-})$ and $\xi_{\rho}^{d-1}(\gamma^{-})$) are the attracting fixed points of $\rho(\gamma)$ (resp. $\rho(\gamma^{-1})$) in $\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ and $\mathsf{Gr}_{d-1}(\operatorname{\mathbb{R}}^{d})$ respectively. Now since $\Lambda_{\Gamma_{0}}$ is infinite we may find $\delta\in\Gamma_{0}$ such that $\{\delta\gamma^{+},\delta\gamma^{-}\}\cap\{\gamma^{+},\gamma^{-}\}$ is empty. Since, up to subsequence, $\lim_{n}\gamma^{\pm n}\delta^{\pm 1}\gamma^{+}=\lim_{n}\gamma^{\pm n}\delta^{\pm 1}\gamma^{-}=\gamma^{\pm}$, we have
@@
This shows that the set $\big{\{}\rho(\gamma),\rho(\gamma^{-1}),\rho(\delta\gamma\delta^{-1}),\rho(\delta\gamma^{-1}\delta^{-1})\big{\}}$ is well positioned in the sense of [CLS17, App. A]. Then by a result of Kapovich–Leeb–Porti [KLP14, Thm. 7.40] and Canary–Lee–Stover–Sambarino [CLS17, Thm. A2], there exists $N>1$ such that the group $\big{\langle}\rho(\gamma^{N}),\rho(\delta\gamma^{N}\delta^{-1})\big{\rangle}$ is a free and $1$-Anosov subgroup of $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$. ∎
@@ 2.6. Relatively Anosov subgroups preserving domains
The following result was established for hyperbolic groups with 1-Anosov representations in [IZ20, Prop. 5.3] (see also [Zim21, Thm. 3.1] and [CT20, Prop. 2.8]), and also holds in the relative case with essentially the same proof.
@@
Let $(\Gamma,\operatorname{\mathcal{P}})$ is a relatively hyperbolic group such that $\partial(\Gamma,\operatorname{\mathcal{P}})$ is connected. Suppose $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{PGL}}_{d}(\operatorname{\mathbb{R}})$ is a relatively 1-Anosov representation, and $\xi_{\rho}^{1}\colon\partial(\Gamma,\operatorname{\mathcal{P}})\to\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ and $(\xi_{\rho}^{1})^{*}\colon\partial(\Gamma,\operatorname{\mathcal{P}})\to\mathbb{P}((\operatorname{\mathbb{R}}^{d})^{*})$ are the associated limit maps.
@@
The proof of (1) follows that of [IZ20, Prop. 5.3(1)].
The proof of (2) follows the arguments in [Zim21] or [CT20, Prop. 2.8].
For the proof of (3), we modify the proof of [IZ20, Prop. 5.3(2)], noting that the result of Swarup used in that argument — if a hyperbolic group has connected Gromov boundary $\partial_{\infty}\Gamma$, then $\partial_{\infty}\Gamma$ has no (global) cut-points — is no longer true in the relative case. Instead, we use Proposition 2.3, which gives a non-cut pair $\{x,y\}$.
Fix an affine chart $\mathbb{A}$ containing $\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}}))$. Working in $\mathbb{P}(\mathrm{span}(\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}})))$, and arguing as in [Zim21, §3.1], we obtain a $\Gamma$-invariant properly convex domain $\Omega_{0}\subset\mathbb{P}(\mathrm{span}(\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}})))$ that $\Omega_{0}$ avoids $\ker((\xi_{\rho}^{1})^{*}(x))$ and $\ker((\xi_{\rho}^{1})^{*}(y))$. Since $\Omega_{0}$ is $\rho(\Gamma)$-invariant, it also avoids all of the $\rho(\Gamma)$-translates of $\ker((\xi_{\rho}^{1})^{*}(x))$ and $\ker((\xi_{\rho}^{1})^{*}(y))$. By minimality of the $\Gamma$-action on $\partial(\Gamma,\operatorname{\mathcal{P}})$ and continuity and equivariance of the limit maps, we conclude that $\Omega_{0}\cap\ker\xi_{\rho}^{*}(z)=\varnothing$ for all $z\in\partial(\Gamma,\operatorname{\mathcal{P}})$.
Then take a bounded neighborhood $N\supset\Omega_{0}$ in our affine chart $\mathbb{A}$. Arguing as in [IZ20, Lem. 5.13], given any $p\in\Omega_{0}$ we may find a neighborhood $U\subset\mathbb{A}$ of $p$ whose $\rho(\Gamma)$-orbit is contained in $N$. Then we may take our convex domain to be the (relative interior of the) convex hull of $\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}}))\cup\rho(\Gamma)\cdot N$. ∎
@@ 3.1. Convergence groups and limit maps
Let $X$ be a compact metrizable space and $\Gamma$ be a finitely-generated group acting as a non-elementary convergence group on $X$. Suppose $\rho\colon\Gamma\to\operatorname{\mathsf{GL}}_{d}(\operatorname{\mathbb{R}})$ is a representation which admits a non-constant continuous $\rho$-equivariant map $\xi\colon X\to\mathbb{P}(\operatorname{\mathbb{R}}^{d})$.
@@
Note that the map $\xi$ is also $\hat{\rho}$-equivariant. If $\hat{\rho}$ is not discrete or has infinite kernel, then there exists an infinite sequence of elements $(\gamma_{n})_{n\in\mathbb{N}}$ of $\Gamma$ such that $\lim_{n}\hat{\rho}(\gamma_{n})=\textup{I}_{d}$. Note that up to passing to a subsequence there exist $\eta_{1},\eta_{2}\in X$ such that $\lim\gamma_{n}x=\eta_{1}$ for every $x\in X\smallsetminus\{\eta_{2}\}$. Then observe that $\xi(x)=\lim_{n}\rho(\gamma_{n})\xi(x)=\lim_{n}\xi(\gamma_{n}x)=\xi(\eta_{1})$ for every $x\in X\smallsetminus\{\eta_{2}\}$. In particular, by continuity, $\xi$ is constant and this is a contradiction. It follows that $\hat{\rho}(\Gamma)$ is a discrete subgroup of $\operatorname{\mathsf{GL}}_{d}(\operatorname{\mathbb{R}})$ and $\hat{\rho}$ has finite kernel. ∎
@@ 3.2. The $\operatorname{\mathsf{SL}}_{3}(\operatorname{\mathbb{R}})$ case
It remains to deal with the case where $\rho$ is non-faithful. In this case, note that $\ker\rho$ is finite. Then the above arguments all apply for the quotients $\Gamma^{\prime}:=\Gamma/\ker\rho$ and $\Gamma^{\prime}_{0}:=\Gamma_{0}/\ker\rho$, and show that one of the following holds: either $\Gamma^{\prime}_{0}$ is virtually nilpotent or virtually free, or Γsuperscriptnormal-$\Gamma^{\prime}$ is virtually a surface group. The conclusion then follows by Proposition 2.4 and Theorem 2.2. ∎
@@ 3.3. Anosov representations into symplectic groups
Let $V_{1}:=\textup{span}(\xi_{\rho}^{1}(\Lambda_{\Gamma_{0}}))$ and observe that $\xi_{\rho}^{1}(\Lambda_{\Gamma_{0}})$ is a connected compact subset of the affine chart $B_{\eta}:=\mathbb{P}(V_{1})\smallsetminus\mathbb{P}(V_{1})\cap\mathbb{P}(\xi_{\rho}^{2m-1}(\eta))$ of $\mathbb{P}(V_{1})$. In particular, $\rho|_{V_{1}}(\Gamma_{0})$ (and hence $\rho|_{V_{1}}(\langle a,b\rangle)$) preserves a properly convex domain $\Omega_{0}\subset\mathbb{P}(V_{1})$. We deduce that there exists a finite-index subgroup $H_{0}\subset\langle a,b\rangle$ such that $\rho|_{V_{1}}(H_{0})$ preserves a properly convex cone $C_{1}\subset V_{1}$ and hence $\rho|_{V_{1}}(H_{0})$ is positively proximal (see [Ben00]).
Since every $h\in H_{0}\smallsetminus\{1\}$ admits an eigenvector of maximum modulus in $V_{1}$ (because the limit set of $\rho(H_{0})$ is contained in $\mathbb{P}(V_{1})$), we deduce that $\lambda_{1}(\rho(h))>0$. Since $\rho|_{H_{0}}\colon H_{0}\to\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ is $1$-Anosov, by the openness of Anosov representations [GW12, Lab06], we may find a continuous path of representations $\left\{\rho_{t}\colon H_{0}\to\operatorname{\mathsf{Sp}}_{2m}(\operatorname{\mathbb{R}})\right\}_{t\in[0,1]}$ such that $\rho_{0}=\rho|_{H_{0}}$, $\rho_{1}$ is Zariski-dense in $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ and $\rho_{t}$ is $1$-Anosov for every $0\leq t\leq 1$. Now for $h\in H_{0}\smallsetminus\{1\}$, $\rho_{t}(h)$ is $1$-proximal, so by continuity we obtain
@@
We deduce that $\rho_{1}(H_{0})$ is a strongly irreducible subgroup of $\operatorname{\mathsf{GL}}_{2m}(\operatorname{\mathbb{R}})$ (since it is a Zariski dense subgroup of $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$) all of whose non-trivial elements are positively $1$-proximal. Thus, by [Ben00], $\rho_{1}(H_{0})$ preserves a properly convex cone $\mathcal{C}^{\prime}\subset\operatorname{\mathbb{R}}^{2m}$. However, this is contradiction by Benoist’s work, see [Ben00, Cor. 3.5].
Therefore, we conclude that $\partial(\Gamma,\operatorname{\mathcal{P}})=\Lambda_{\Gamma_{0}}$. In fact, our arguments show that $\xi_{\rho}^{1}(\Lambda_{L}))$ is not bounded in any affine chart of $\mathbb{P}(\operatorname{\mathbb{R}}^{2m})$, as soon as $\Lambda_{L}\subset\Gamma$ is connected and $L$ is a non-peripheral subgroup of $\Gamma$.
@@ 3.4. The case of the disconnected boundary
In the particular case where of a relatively $1$-Anosov representation $\rho\colon(\Gamma,\mathcal{P})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$, by [DT23, Cor. 1.5], $\Gamma$ is Gromov hyperbolic, every nilpotent vertex group of $\mathcal{G}$ is virtually cyclic, and hence $\Gamma$ is virtually free. Therefore, if $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\mathsf{SL}_{3}(\mathbb{R})$ is relatively Anosov and $\partial(\Gamma,\operatorname{\mathcal{P}})$ is disconnected, then $\Gamma$ is virtually free.∎
@@
As in the proof of Theorem 3.2, the bulk of the argument involves looking at what happens in the case where $\Lambda_{\Gamma_{0}}$ is neither a singleton nor a circle, and hence we are not in cases (i) or (ii). To this end, we will establish a few lemmas about subgroups of $\Gamma$ with connected limit set in $\partial(\Gamma,\operatorname{\mathcal{P}})$ not a point and not homeomorphic to a circle. We first prove the following result, which may be seen as a relatively easier case of [HW22, Thm. 8.1]. We note that all of our groups and peripheral subgroups are finitely generated.
@@
By our hypotheses, the parabolic action of each $P\in\operatorname{\mathcal{P}}$ on $\xi(\partial(\Gamma,\operatorname{\mathcal{P}}))$ extends to a parabolic action of $P$ on all of the 2-sphere $S$. Hence $P$ acts properly on the plane $S\smallsetminus\{x_{P}\}$, where $x_{P}$ denotes the fixed point of $P$, and therefore $P$ is virtually a surface group or free group (see [HW22, Cor. 3.2]). ∎
@@
We might hope to conclude that $\Gamma$ is a Kleinian group. This could follow from an extension to the Cannon conjecture, which remains a major open question. (See [HW22, Conj. 1.1] for a statement given the additional assumption that $\partial(\Gamma,\operatorname{\mathcal{P}})$ has no cut points.) In our case, we can at least obtain that $\Gamma$ is the fundamental group of a 3-manifold:
@@
We may assume $M$ is compact by the Scott core theorem [Sco73]. This states that: if $\Gamma$ is the fundamental group of a 3-manifold $M$ (possibly with boundary) and $\Gamma$ is finitely-generated, then there exists a compact three-dimensional submanifold $N\subset M$, called the compact core, such that the inclusion $\iota\colon N\hookrightarrow M$ induces an isomorphism on fundamental groups. ∎
@@
If $\Gamma_{0}\cong\pi_{1}N^{3}$ where $N^{3}$ is a compact 3-manifold admitting a JSJ decomposition (this happens e.g. when $N^{3}$ has empty or toroidal boundary, see [AFW15, Thm. 7.2.2]), then we can further conclude, using the JSJ decomposition, that $N^{3}$ consists of hyperbolizable pieces joined across tori.

### v2

@@ 2.5. Relatively Anosov subgroups
Let $(\Gamma,\mathcal{P})$ be a non-elementary relatively hyperbolic group which is virtually torsion-free. The limit set of a subgroup $\Gamma^{\prime}$ of $\Gamma$ is a singleton if and only if $\Gamma^{\prime}$ is virtually contained in a conjugate of a peripheral subgroup.
@@
By passing to exterior powers we may assume that $k=1$, see [reldomreps, Def. 10.2] and [ZZ1, Ex. 13.2]). Since $\Gamma_{0}$ is non-elementary there exists $\gamma\in\Gamma$ such that $\rho(\gamma)$ is $1$-biproximal (see [reldomreps, Prop. 4.6] and [ZZ1, Prop. 4.2]), and $\xi_{\rho}^{1}(\gamma^{+})$ and $\xi_{\rho}^{d-1}(\gamma^{+})$ (resp. $\xi_{\rho}^{1}(\gamma^{-})$ and $\xi_{\rho}^{d-1}(\gamma^{-})$) are the attracting fixed points of $\rho(\gamma)$ (resp. $\rho(\gamma^{-1})$) in $\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ and $\mathsf{Gr}_{d-1}(\operatorname{\mathbb{R}}^{d})$ respectively. Now since $\Lambda_{\Gamma_{0}}$ is infinite we may find $\delta\in\Gamma_{0}$ such that $\{\delta\gamma^{+},\delta\gamma^{-}\}\cap\{\gamma^{+},\gamma^{-}\}$ is empty. Since, up to subsequence, $\lim_{n}\gamma^{\pm n}\delta^{\pm 1}\gamma^{+}=\lim_{n}\gamma^{\pm n}\delta^{\pm 1}\gamma^{-}=\gamma^{\pm}$, we have
@@
This shows that the set $\big{\{}\rho(\gamma),\rho(\gamma^{-1}),\rho(\delta\gamma\delta^{-1}),\rho(\delta\gamma^{-1}\delta^{-1})\big{\}}$ is well positioned in the sense of [CLS, App. A]. Then by a result of Kapovich–Leeb–Porti [KLP, Thm. 7.40] and Canary–Lee–Stover–Sambarino [CLS, Thm. A2], there exists $N>1$ such that the group $\big{\langle}\rho(\gamma^{N}),\rho(\delta\gamma^{N}\delta^{-1})\big{\rangle}$ is a free and $1$-Anosov subgroup of $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$. ∎
@@ 2.6. Relatively Anosov subgroups preserving domains
The following result was established for hyperbolic groups with 1-Anosov representations in [IZ, Prop. 5.3] (see also [Zimmer, Thm. 3.1] and [CanaryTsouvalas, Prop. 2.8]), and also holds in the relative case with essentially the same proof.
@@
Let $(\Gamma,\operatorname{\mathcal{P}})$ is a relatively hyperbolic group such that $\partial(\Gamma,\operatorname{\mathcal{P}})$ is connected. Suppose $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{PGL}}_{d}(\operatorname{\mathbb{R}})$ is a relatively 1-Anosov representation, and
$\left(\xi_{\rho}^{1},(\xi_{\rho}^{1})^{*}\right)\colon\partial(\Gamma,\operatorname{\mathcal{P}})\to\mathbb{P}(\operatorname{\mathbb{R}}^{d})\times\mathbb{P}((\operatorname{\mathbb{R}}^{d})^{*})$
are the associated limit maps.
@@
The proof of (1) follows that of [IZ, Prop. 5.3(1)].
The proof of (2) follows the arguments in [Zimmer] or [CanaryTsouvalas, Prop. 2.8].
For the proof of (3), we modify the proof of [IZ, Prop. 5.3(2)], noting that the result of Swarup used in that argument — if a hyperbolic group has connected Gromov boundary $\partial_{\infty}\Gamma$, then $\partial_{\infty}\Gamma$ has no (global) cut-points — is no longer true in the relative case. Instead, we use Proposition 2.3, which gives a non-cut pair $\{x,y\}$.
Fix an affine chart $\mathbb{A}$ containing $\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}}))$. Working in $\mathbb{P}(\mathrm{span}(\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}})))$, and arguing as in [Zimmer, §3.1], we obtain a $\Gamma$-invariant properly convex domain $\Omega_{0}\subset\mathbb{P}(\mathrm{span}(\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}})))$ that $\Omega_{0}$ avoids $\ker((\xi_{\rho}^{1})^{*}(x))$ and $\ker((\xi_{\rho}^{1})^{*}(y))$. Since $\Omega_{0}$ is $\rho(\Gamma)$-invariant, it also avoids all of the $\rho(\Gamma)$-translates of $\ker((\xi_{\rho}^{1})^{*}(x))$ and $\ker((\xi_{\rho}^{1})^{*}(y))$. By minimality of the $\Gamma$-action on $\partial(\Gamma,\operatorname{\mathcal{P}})$ and continuity and equivariance of the limit maps, we conclude that $\Omega_{0}\cap\ker\xi_{\rho}^{*}(z)=\varnothing$ for all $z\in\partial(\Gamma,\operatorname{\mathcal{P}})$.
Then take a bounded neighborhood $N\supset\Omega_{0}$ in our affine chart $\mathbb{A}$. Arguing as in [IZ, Lem. 5.13], given any $p\in\Omega_{0}$ we may find a neighborhood $U\subset\mathbb{A}$ of $p$ whose $\rho(\Gamma)$-orbit is contained in $N$. Then we may take our convex domain to be the (relative interior of the) convex hull of $\xi_{\rho}^{1}(\partial(\Gamma,\operatorname{\mathcal{P}}))\cup\rho(\Gamma)\cdot N$. ∎
@@ 3.1. Convergence groups and limit maps
Let $X$ be a compact metrizable space and $\Gamma$ be a finitely-generated group acting (minimally) as a non-elementary convergence group on $X$. Suppose $\rho\colon\Gamma\to\operatorname{\mathsf{GL}}_{d}(\operatorname{\mathbb{R}})$ is a representation which admits a non-constant continuous $\rho$-equivariant map $\xi\colon X\to\mathbb{P}(\operatorname{\mathbb{R}}^{d})$.
@@
Note that the map $\xi$ is also $\hat{\rho}$-equivariant. If $\hat{\rho}$ is not discrete or has infinite kernel, then there exists an infinite sequence of elements $(\gamma_{n})_{n\in\mathbb{N}}$ of $\Gamma$ such that $\lim_{n}\hat{\rho}(\gamma_{n})=\textup{I}_{d}$. Note that up to passing to a subsequence there exist $\eta_{1},\eta_{2}\in X$ such that $\lim\gamma_{n}x=\eta_{1}$ for every $x\in X\smallsetminus\{\eta_{2}\}$. Then observe that $\xi(x)=\lim_{n}\rho(\gamma_{n})\xi(x)=\lim_{n}\xi(\gamma_{n}x)=\xi(\eta_{1})$ for every $x\in X\smallsetminus\{\eta_{2}\}$. In particular, by continuity, and because $X$ is perfect and hence $\eta_{2}$ is not isolated, $\xi$ is constant and this is a contradiction. It follows that $\hat{\rho}(\Gamma)$ is a discrete subgroup of $\operatorname{\mathsf{GL}}_{d}(\operatorname{\mathbb{R}})$ and $\hat{\rho}$ has finite kernel. ∎
@@ 3.2. The $\operatorname{\mathsf{SL}}_{3}(\operatorname{\mathbb{R}})$ case
It remains to deal with the case where $\rho$ is non-faithful. In this case, note that $\ker\rho$ is finite. Then the above arguments all apply for the quotients $\Gamma^{\prime}:=\Gamma/\ker\rho$ and $\Gamma^{\prime}_{0}:=\Gamma_{0}/\ker\rho$, and show that one of the following holds: either $\Gamma^{\prime}_{0}$ is virtually nilpotent or virtually free, or $\Gamma^{\prime}$ is virtually a surface group. The conclusion then follows by Proposition 2.4 and Theorem 2.2. ∎
@@ 3.3. Anosov representations into symplectic groups
Let $V_{1}:=\textup{span}(\xi_{\rho}^{1}(\Lambda_{\Gamma_{0}}))$ and observe that $\xi_{\rho}^{1}(\Lambda_{\Gamma_{0}})$ is a connected compact subset of the affine chart $B_{\eta}:=\mathbb{P}(V_{1})\smallsetminus\mathbb{P}(V_{1})\cap\mathbb{P}(\xi_{\rho}^{2m-1}(\eta))$ of $\mathbb{P}(V_{1})$. In particular, $\rho|_{V_{1}}(\Gamma_{0})$ (and hence $\rho|_{V_{1}}(\langle a,b\rangle)$) preserves a properly convex domain $\Omega_{0}\subset\mathbb{P}(V_{1})$. We deduce that there exists a finite-index subgroup $H_{0}\subset\langle a,b\rangle$ such that $\rho|_{V_{1}}(H_{0})$ preserves a properly convex cone $C_{1}\subset V_{1}$ and hence $\rho|_{V_{1}}(H_{0})$ is positively proximal (see [Ben]).
Since every $h\in H_{0}\smallsetminus\{1\}$ admits an eigenvector of maximum modulus in $V_{1}$ (because the limit set of $\rho(H_{0})$ is contained in $\mathbb{P}(V_{1})$), we deduce that $\lambda_{1}(\rho(h))>0$. Since $\rho|_{H_{0}}\colon H_{0}\to\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ is $1$-Anosov, by the openness of Anosov representations [GW, Lab], we may find a continuous path of representations $\left\{\rho_{t}\colon H_{0}\to\operatorname{\mathsf{Sp}}_{2m}(\operatorname{\mathbb{R}})\right\}_{t\in[0,1]}$ such that $\rho_{0}=\rho|_{H_{0}}$, $\rho_{1}$ is Zariski-dense in $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ and $\rho_{t}$ is $1$-Anosov for every $0\leq t\leq 1$. Now for $h\in H_{0}\smallsetminus\{1\}$, $\rho_{t}(h)$ is $1$-proximal, so by continuity we obtain
@@
We deduce that $\rho_{1}(H_{0})$ is a strongly irreducible subgroup of $\operatorname{\mathsf{GL}}_{2m}(\operatorname{\mathbb{R}})$ (since it is a Zariski dense subgroup of $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$) all of whose non-trivial elements are positively $1$-proximal. Thus, by [Ben], $\rho_{1}(H_{0})$ preserves a properly convex cone $\mathcal{C}^{\prime}\subset\operatorname{\mathbb{R}}^{2m}$. However, this is contradiction by Benoist’s work, see [Ben, Cor. 3.5].
Therefore, we conclude that $\partial(\Gamma,\operatorname{\mathcal{P}})=\Lambda_{\Gamma_{0}}$. In fact, our arguments show that $\xi_{\rho}^{1}(\Lambda_{L}))$ is not bounded in any affine chart of $\mathbb{P}(\operatorname{\mathbb{R}}^{2m})$, as soon as $\Lambda_{L}$ is connected and $L$ is not contained in a conjugate of a peripheral subgroup of $\Gamma$.
@@ 3.4. The case of the disconnected boundary
In the particular case where of a relatively $1$-Anosov representation $\rho\colon(\Gamma,\mathcal{P})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$, by [DoubaT, Cor. 1.5], $\Gamma$ is Gromov hyperbolic, every nilpotent vertex group of $\mathcal{G}$ is virtually cyclic, and hence $\Gamma$ is virtually free. Therefore, if $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\mathsf{SL}_{3}(\mathbb{R})$ is relatively Anosov and $\partial(\Gamma,\operatorname{\mathcal{P}})$ is disconnected, then $\Gamma$ is virtually free.∎
@@
As in the proof of Theorem 3.2, the bulk of the argument involves looking at what happens in the case where $\Lambda_{\Gamma_{0}}$ is neither a singleton nor a circle, and hence we are not in cases (i) or (ii). To this end, we will establish a few lemmas about subgroups of $\Gamma$ with connected limit set in $\partial(\Gamma,\operatorname{\mathcal{P}})$ not a point and not homeomorphic to a circle. We first prove the following result, which may be seen as a relatively easier case of [HruskaWalsh, Thm. 8.1]. We note that all of our groups and peripheral subgroups are finitely generated.
@@
By our hypotheses, the parabolic action of each $P\in\operatorname{\mathcal{P}}$ on $\xi(\partial(\Gamma,\operatorname{\mathcal{P}}))$ extends to a parabolic action of $P$ on all of the 2-sphere $S$. Hence $P$ acts properly on the plane $S\smallsetminus\{x_{P}\}$, where $x_{P}$ denotes the fixed point of $P$, and therefore $P$ is virtually a surface group or free group (see [HruskaWalsh, Cor. 3.2]). ∎
@@
We might hope to conclude that $\Gamma$ is a Kleinian group. This could follow from an extension to the Cannon conjecture, which remains a major open question. (See [HruskaWalsh, Conj. 1.1] for a statement given the additional assumption that $\partial(\Gamma,\operatorname{\mathcal{P}})$ has no cut points.) In our case, we can at least obtain that $\Gamma$ is the fundamental group of a 3-manifold:
@@
We may assume $M$ is compact by the Scott core theorem [Scott]. This states that: if $\Gamma$ is the fundamental group of a 3-manifold $M$ (possibly with boundary) and $\Gamma$ is finitely-generated, then there exists a compact three-dimensional submanifold $N\subset M$, called the compact core, such that the inclusion $\iota\colon N\hookrightarrow M$ induces an isomorphism on fundamental groups. ∎
@@
If $\Gamma_{0}\cong\pi_{1}N^{3}$ where $N^{3}$ is a compact 3-manifold admitting a JSJ decomposition (this happens e.g. when $N^{3}$ has empty or toroidal boundary, see [AFW, Thm. 7.2.2]), then we can further conclude, using the JSJ decomposition, that $N^{3}$ consists of hyperbolizable pieces joined across tori.